# Evaluating and Mitigating LLM-as-a-judge Bias in Communication Systems

### v1

@@ IV-A Robust Prompt Design and Reasoning
Carefully crafting the evaluation prompt can preempt many biases. Prompts should include explicit instructions that guide the judge model to focus on factual correctness and relevance, and to ignore irrelevant attributes such as the content’s author identity or stylistic flair. For example, the prompt might say “Disregard any gender or identity information and evaluate only on answer quality”. Additionally, we can employ advanced reasoning strategies like chain-of-thought (CoT) prompting to encourage the model to work through a step-by-step, logical evaluation. This structured reasoning helps the judge to base its scores on substantive content rather than superficial features, which is crucial in communication domains (e.g., ensuring a chatbot’s concise but correct reply isn’t undervalued compared to a verbose one).
@@ IV-C Model Calibration and Specialized Training
Calibration techniques can also be applied to down-weight superficial qualities in scoring. For proprietary judges, this might involve probability- or prompt-level calibration that subtracts points for factors like excessive verbosity or unwarranted confidence. The goal is to ensure the scoring criteria align with true content quality rather than presentation. For open-source judges, we can employ bias-focused fine-tuning, that is training the model with curated examples that include “trap” scenarios. In such training, negative samples are provided where an answer looks polished (fluent, lengthy) but violates instructions, and the model is explicitly taught to score these lower. This helps the judge learn to favor instruction-following and correctness over surface shine. Moreover, fine-tuning on domain-specific data (e.g., communication-related QA or moderation examples) further aligns the judge with the contextual norms of that domain, enabling it to better discern what constitutes fairness and relevance [15].
@@ V Conclusion
We evaluated 2 leading LLM-as-a-Judge models on 11 bias types across 4 test datasets and outlined 4 potential mitigation strategies to reduce the bias impact. By shedding light on these judgment biases and their possible remedies, this study aims to promote the development of fairer and more trustworthy LLM-as-a-Judge systems for future communication services.

### v3

@@ IV-A Robust Prompt Design and Reasoning
Carefully crafting the evaluation prompt can preempt many biases. Prompts should include explicit instructions that guide the judge model to focus on factual correctness and relevance, and to ignore irrelevant attributes such as the content’s author identity or stylistic flair [6]. For example, the prompt might say “Disregard any gender or identity information and evaluate only on answer quality”. Additionally, we can employ advanced reasoning strategies like chain-of-thought (CoT) prompting to encourage the model to work through a step-by-step, logical evaluation. This structured reasoning helps the judge to base its scores on substantive content rather than superficial features, which is crucial in communication domains (e.g., ensuring a chatbot’s concise but correct reply isn’t undervalued compared to a verbose one).
@@ IV-C Model Calibration and Specialized Training
Calibration techniques can also be applied to down-weight superficial qualities in scoring. For proprietary judges, this might involve probability- or prompt-level calibration that subtracts points for factors like excessive verbosity or unwarranted confidence. The goal is to ensure the scoring criteria align with true content quality rather than presentation. For open-source judges, we can employ bias-focused fine-tuning, that is training the model with curated examples that include “trap” scenarios. In such training, negative samples are provided where an answer looks polished (fluent, lengthy) but violates instructions, and the model is explicitly taught to score these lower [6]. This helps the judge learn to prioritize instruction-following and correctness over superficial stylistic features. Moreover, fine-tuning on domain-specific data (e.g., communication-related QA or moderation examples) further aligns the judge with the contextual norms of that domain, enabling it to better discern what constitutes fairness and relevance.
@@ V Conclusion
We evaluated 6 leading LLM-as-a-Judge models on 11 bias types across 4 test datasets and outlined 4 potential mitigation strategies to reduce the bias impact. By shedding light on these judgment biases and their possible remedies, this study aims to promote the development of fairer and more trustworthy LLM-as-a-Judge systems for future communication services.